Range and domain of sinx*cosx

Grade:12

1 Answers

Rinkoo Gupta
9 years ago
Domain is R
Range
sinxcosx=2sinxcosx/2=sin2x/2
-1=<sin2x<=1
-1/2=<sin2x/2<=1/2
so Range =[-1/2,1/2]
Thanks & Regards
Rinkoo Gupta
AskIITians faculty